OpenStudy (anonymous):

How would you write 0.3y + 1.2 = –0.4x in standard form?

OpenStudy (anonymous):

I guess it would be 0.4x+0.3y = -1.2

OpenStudy (anonymous):

Usually want y = mx + b form get rid of 0.3 multiplying y by dividing all by 0.3. Move x term to right-hand side.

OpenStudy (whpalmer4):

Standard form is \[Ax+By=C\] with \(A>0\) and preferably \(A, B\) relatively prime integers.

OpenStudy (whpalmer4):

So I would multiply the whole thing by 10 to give \[4x+3y=-12\]